In a shop, the price of a monthly magazine is $m and the price of a weekly magazine is $(m-0.75). One day, the shop receives $168 from selling monthly magazines and $207 from selling weekly magazines. The total number of these magazines sold during this day is 100. Calculate the number of weekly and monthly magazines sold.
Answer by mananth(16946)

In a shop, the price of a monthly magazine is $m and the price of a weekly magazine is $(m-0.75). One day, the shop receives $168 from selling monthly magazines and $207 from selling weekly magazines. The total number of these magazines sold during this day is 100. Calculate the number of weekly and monthly magazines sold.

the price of a monthly magazine is $m
the shop receives $168 from selling monthly magazines
Therefore number of monthly magzines sold is 168/m
the price of a weekly magazine is $(m-0.75)
$207 from selling weekly magazines.
Therefore number of weekly magzines sold is 207/(m-0.75)
The total number of these magazines sold during this day is 100.
168/m +207/(m-0.75)=100
Multiply the equation m(m-0.75)
168*(m-0.75) + 207*m= 100m(m-0.75)
168m-126+207m= 100m^2-75m
375m-126=100m^2-75m
100m^2-450m+126=0
solve
m= 4.2 or 0.3
0.3 not possible
m=4.20
Therefore number of monthly magzines sold is 168/4.20=40


(m-0.75) = 3.45
Therefore number of weekly magzines sold is 207/3.45 =60